OpenStudy (anonymous):

Would this problem be graphing, substitution, or elimination? A perfume maker has stocks of two perfumes on hand. Perfume A sells for $15 per ounce. Perfume B sells for $35 per ounce. How much of each should be combined to make a 3-oz bottle of perfume that can be sold for $63?

OpenStudy (anonymous):

let a be the amt of perfume A, and b be the amt of perfume B. we have been given that \[a+b = 3\] and that \[a*15 + b*35 = 63\] . From equation 1, we substitute \[ b = 3 -a \] so as to get, \[a*15 + (3-a)*35 = 63\] which means, \[15a + 105 - 35a = 63\] which means, \[-20a = -42\] so a = 2.1 ounce and b = 0.9 ounce